30. Prove that|AU BI+AUC + BUC|S|A|+BI+ CI+ AUBUC for any three finite sets A, B and C. DIIDID DODATION
Answer #1

Note that LAUBL - JA1+ B] - LANB HAUCKLARTICL_- lanc) -T-Burt=1B1+ch--BAC)- _And_LA_U.B.UCI = TAI +113] +16) - TADB) :-LANCIB